OpenStudy (anonymous):

The length of a room is double its breadth. Its height is 3m. The area of four walls excluding a door of dimension 4m x 2m is 100 meter square. Find its volume

OpenStudy (anonymous):

breadth=x L=2x H=3

hartnn (hartnn):

Area of 4 walls -4*2 =100 Area of 4 walls = 108 2 (L*H + breadth *H) = 108 2 (6x+3x)=108 can you find 'x' now ?

hartnn (hartnn):

also, did you get those steps ? need any explanation of any step ?

OpenStudy (anonymous):

how 108 cmes

OpenStudy (anonymous):

wait i got it

OpenStudy (anonymous):

now, volume =

hartnn (hartnn):

area of four walls excluding a door of dimension 4m x 2m = Area -4*2 is 100 meter square. =100 ok

hartnn (hartnn):

volume = L*H*breadth = x*2x* 3

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

6*12*3=216

hartnn (hartnn):

yes, thats what i also get.

OpenStudy (anonymous):

@hartnn thank u so much